OpenStudy (anonymous):

t squared plus 6 t over the quantity t minus 1 end quantity minus 7 over quantity t minus 1

OpenStudy (anonymous):

\[\frac{ t^2+6t }{ (t-1) }-\frac{ 7 }{ (t-1) }\]Is this what you are asking about?

OpenStudy (anonymous):

Well if it does happen to be the problem, this is how you would solve this. You have a commond denominator, so simply subtract and combine, factor the polynomial, cancle like terms, and then you'll have your t+7 as your answer.\[=\frac{ t^2+6t-7 }{ (t-1) }=\frac{ (t+7)(t-1) }{ (t-1) }= t+7\]

OpenStudy (anonymous):

And i apologize ahead of time for all the grammatical errors in that paragraph, my typing is lagging right now xD